OpenStudy (anonymous):

A copy machine makes 60 papers per minute a second copy machine makes 80 copies per minute. The second machine starts making copies 2 minutes after the first machine starts. Both machines stop making copies 8 minutes after the first machine started. Together, the two machines made how many copies?

OpenStudy (anonymous):

1st machine makes copies for 8 minutes, so the 2nd makes copies for 8 - 2 = 6 minutes. 1. How many copies can the 1st machine make in 8 minutes? 2. How many copies can the 2nd machine make in 6 minutes?

OpenStudy (anonymous):

60*8=? 80*6=?

OpenStudy (anonymous):

yes, then add them together

OpenStudy (anonymous):

what do you think it is @aldaze412 .

OpenStudy (anonymous):

i was getting there lol

OpenStudy (anonymous):

960!!! Is that right? Thanks guys :D I was solving it a totally different way 😔

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

Thank you ! 😁

OpenStudy (anonymous):

agreed that is correct

OpenStudy (anonymous):

np bud